OpenStudy (jmartinez638):

The surface area of cube is growing at a constant rate of 54 ft^2/min. What is the rate of change of the volume of the cube when the cube has side length 20?

OpenStudy (jack1):

do u understand (mathematically) the relationship between volume and surface area?

OpenStudy (irishboy123):

\(A = 6 x^2 \\ \dot A = 12 x \; \dot x\) \(V = x^3\) \(\dot V = 3 x^2 \; \dot x = \dots\)

OpenStudy (jmartinez638):

*A little confused*

OpenStudy (irishboy123):

are you learning/applying calculus?

OpenStudy (jmartinez638):

I am learning calculus, but I'm not so so far into it. Just touching related rates and need help working through this problem

OpenStudy (irishboy123):

the chain rule seems important here. you know that? \(\dfrac{dy}{dt} = \dfrac{dy}{dx}\dfrac{dx}{dt}\)

OpenStudy (jmartinez638):

yes

OpenStudy (irishboy123):

and \(A = 6 x^2 \) should make great sense too as \(\dfrac{ dA}{dt} = \dfrac{dA}{dx} \dfrac{dx}{dt} = 12 x \; \dfrac{ dx}{dt} \)

OpenStudy (irishboy123):

"as" :-(( so

OpenStudy (jmartinez638):

Ah, that makes sense. Thanks